the standard enthalpy of combustion of ethane and hydrogen are -1559.7 kJ/mol and -286kJ/mol respectively, while the standard enthalpy of hydrogenation of ethyne to ethane is -296kJ/mol.
Use Hess'ss law to find the enthalpy of combustion of ethyne
1
Expert's answer
2019-12-11T07:54:32-0500

ΔHcomb(C2H2)=ΔHcomb(C2H6)-2ΔHcomb(H2)+ΔHhydr(H2)=-1559.7+(-296)-2(-286)=-1283.7 kJ/mol